OpenStudy (anonymous):

A square pyramid has a slant height of 13 meters and a height of 12 meters. What is the surface area if the square base has sides of length 10 meters?

OpenStudy (anonymous):

I started by linfing the lateral area.. La= 1/2 pl and got 1/2 40(13) = 260

OpenStudy (ash2326):

Let's find the area of the square base= 10 *10=100 m^2 Now let's find the area of one triangular surface its lateral height=13 meters base=10 so area =13*10/2=65 there are 4 surfaces like this=65*4=260 total surface area=260+100=360 m^2

OpenStudy (mertsj):